OpenStudy (anonymous):

What is the value of y? 6(5 + 3) = 6 • 5 + 6 • y

OpenStudy (phebe):

WELCOME TO OPENSTUDY!!!

OpenStudy (anonymous):

y=1

OpenStudy (phebe):

WRONG ITS y = 3

OpenStudy (anonymous):

nope i did the work and it came out y=1

OpenStudy (anonymous):

i got 1 to

OpenStudy (anonymous):

yeah

OpenStudy (phebe):

OH IG IM WRONG THEN LOL SORRY

OpenStudy (anonymous):

y=3

OpenStudy (anonymous):

its alright :)

OpenStudy (anonymous):

thanksss guys :)

OpenStudy (phebe):

SO I WAS RIGHT?!?!?!?!

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

np @msammonss

OpenStudy (phebe):

SO THE ANSWER 3 RIGHT?!?!?!

OpenStudy (♪chibiterasu):

I got 3.

OpenStudy (phebe):

SO I WAS RIGHT THANKS GUYZ!

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so is phebe right?

OpenStudy (♪chibiterasu):

6(5 + 3) = 6(8) = 48 6 * 5 + 6y 30 + 6y 48 - 30 = 18 18/6 = 3

OpenStudy (phebe):

Yes im right! @skyvibe

OpenStudy (♪chibiterasu):

I have no idea where the negatives came from when you were solving but alright.

OpenStudy (phebe):

i agree^^

OpenStudy (anonymous):

alright sorry i probably got confused :)